OpenStudy (anonymous):

the length of a rectangle is 2cm less than 7 times the width. the perimeter is 60. find the length and width.

OpenStudy (anonymous):

where are you stuck ?

OpenStudy (anonymous):

the equation

OpenStudy (asnaseer):

call the width 'w', then: 7 times the width = 7w 2cm less than 7 times the width = 7w-2 so width = w length = 7w-2

OpenStudy (anonymous):

length is 7x-2 width is x 2(x)+2(7x-2) 2x+14x-4=60 16x=60+4 16x=64 x=4 which is width now for length which is 7(4)-2=26